OpenStudy (anonymous):

given the linear equation x+y=5, find the slope of its graph. A. -5 B. 1 C. -1 D. 5

OpenStudy (arabpride):

into y = mx + b form x+y=5 -x -x y = -x + 5 b = slope in this case your b is the 5 and it is positive, therefore D. 5

OpenStudy (anonymous):

thank you very much

OpenStudy (arabpride):

Welcome *^.^*